evaluating profitability Last​ year, Stevens Inc. had sales of

​$401 comma 000

​,

with a cost of goods sold of

​$111 comma 000

.

The​ firm's operating expenses were

$ 131 comma 000

​,

and its increase in retained earnings was

​$58 comma 000

.

There are currently

22 comma 500

common stock shares outstanding and the firm pays a

​$1.57

dividend per share.

a. Assuming the​ firm's earnings are taxed at

34

​percent, construct the​ firm's income statement.

b. Compute the​ firm's operating profit margin.

c. What was the times interest​ earned?
Answer #1

total dividend=22500*1.57=35325

Net income=increase in retained earnings+total dividend

=58000+35325

=93325

Income statement
Sales        4,01,000
Cost of goods sold       -1,11,000
Gross profit        2,90,000
Operating expenses       -1,31,000
Operating profit        1,59,000
Interest expense          -17,598
Profit before taxes        1,41,402
Taxes          -48,077
Net income            93,325

operating profit margin=159000/401000=39.65%

the times interest​ earned=159000/17598=9.04

the above is answer..